A stone is thrown into a well 6m deep. The stone is thrown from rest. How


long will it take before you hear the stone hit the floor of the well. Assume


speed of sound is 2m/s (this is not the actual speed of sound, please!!)?

1
Expert's answer
2022-03-23T07:15:32-0400

The total time that elapsed as you thrown a stone and you hear the stone hit the floor of the well

"t=t_1+t_2=\sqrt{2h/g}+h/v_s\\
=\sqrt{2*6/9.8}+6/2=4.1\;\rm s"